OpenStudy (anonymous):

If a force of 12 N is applied to an object with a mass of 2 kg, the object will accelerate at A. 0.17 m/s2. B. 24 m/s2. C. 6 m/s2. D. 12 m/s2.

OpenStudy (theeric):

F = m * a is the basic, simple formula here. F is force on the object m is mass of the object a is acceleration of the object F / m = a is the algebraically manipulated formula. From the basic formula up there, divide each side by "m".

OpenStudy (anonymous):

so (B)

OpenStudy (anonymous):

no c

OpenStudy (anonymous):

what eric said is correct

OpenStudy (anonymous):

so how that is how you got (C)

OpenStudy (anonymous):

F=m*a F=12N M=2kg simple algebra and you come up with Erics eq. F/m=a 12N/2kg=6m/s^2

OpenStudy (theeric):

Thank you for the credit. I agree with C.\[6 m/s ^{2}\] Good luck!